# InstructSAM: A Training-Free Framework for Instruction-Oriented Remote Sensing Object Recognition

### v1

@@ Abstract
Language-Guided object recognition in remote sensing imagery is crucial for large-scale mapping and automated data annotation. However, existing open-vocabulary and visual grounding methods rely on explicit category cues, limiting their ability to handle complex or implicit queries that require advanced reasoning. To address this issue, we introduce a new suite of tasks, including Instruction-Oriented Object Counting, Detection, and Segmentation (InstructCDS), covering open-vocabulary, open-ended, and open-subclass scenarios. We further present EarthInstruct, the first InstructCDS benchmark for earth observation. It is constructed from two diverse remote sensing datasets with varying spatial resolutions and annotation rules across 20 categories, necessitating models to interpret dataset-specific instructions. Given the scarcity of semantically rich labeled data in remote sensing, we propose InstructSAM, a training-free framework for instruction-driven object recognition. InstructSAM leverages large vision-language models to interpret user instructions and estimate object counts, employs SAM2 for mask proposal, and formulates mask-label assignment as a binary integer programming problem. By integrating semantic similarity with counting constraints, InstructSAM efficiently assigns categories to predicted masks without relying on confidence thresholds. Experiments demonstrate that InstructSAM matches or surpasses specialized baselines across multiple tasks while maintaining near-constant inference time regardless of object count, reducing output tokens by 89% and overall runtime by over 32% compared to direct generation approaches. We believe the contributions of the proposed tasks, benchmark, and effective approach will advance future research in developing versatile object recognition systems.
@@ 1 Introduction
Recent advancements in VLMs have demonstrated impressive performance in object detection liu_grounding_2023 ; lin2024generative ; pi2024ins , semantic segmentation yu2023convolutions ; cho2024cat , visual grounding bai2025qwen2 ; chen2024expanding and reasoning-based segmentation  ren2024pixellm ; lai2024lisa within the natural image domain. However, transferring these methods to remote sensing imagery presents several challenges. First, direct inference leads to significant accuracy degradation due to the domain gap between natural and aerial images fu2024cross . Second, most existing remote sensing open-vocabulary detection li2024toward ; li2024exploiting ; zang2024zero and segmentation zermatten2025learning ; ye2025towards ; huang2025zori methods are trained on datasets with only a limited number of categories, restricting their generalization to diverse unseen categories. Third, conventional detectors li2022grounded ; cheng2024yolo ; minderer2023scaling rely on a threshold to filter predicted bounding boxes, which is not obtainable in zero-shot scenarios.
@@ 2.2 Instruction-Oriented Remote Sensing Object Detection and Segmentation
Recent advances in VLMs radford2021learning ; liu2023visual have also enabled open-vocabulary learning in the remote sensing domain. Diverse semantic tags from OpenStreetMap haklay2008openstreetmap and those generated by LVLMs drive the development of contrastive language-image pre-training in remote sensing images zhang2024rs5m ; wang2024skyscript . Following the generic open-vocabulary learning frameworks, methods for remote sensing open-vocabulary detection li2024toward ; wei2024ova ; pan2025locate ; huang2025openrsd and segmentation zermatten2025learning ; ye2025towards ; huang2025zori ; li2024segearth methods emerge. However, their human-annotated training data, being limited to a few dozen categories li2020object ; waqas2019isaid ; sun2022fair1m , hinder generalization to out-of-distribution or zero-shot scenarios. Although some LVLMs could support remote sensing tasks like single-class object detection kuckreja2024geochat ; luo2024skysensegpt ; irvin2024teochat , visual grounding muhtar2024lhrs ; pang2025vhm , referring expression segmentation zhou2024geoground ; ou2025geopix , grounded conversation generation shabbir2025geopixel , and scene graph generation luo2024skysensegpt , they failed to follow complex reasoning instructions, such as open-vocabulary and open-subclass object detection. To annotate vast-vocabulary training data for remote sensing object detection, LAE-Label pan2025locate employs a generic LVLM chen2024far to predict categories for cropped mask proposals. However, this approach loses global context for accurate category classification.
@@ 4.3 Mask-Label Matching with Counting Constraints
As depicted in Figure 3, this BIP formulation elegantly fuses visual information, semantic information, and quantitative information. The visual information comes from the CLIP embeddings of mask proposals $\{v_{i}\}$, which contribute to $s_{ij}$. Semantic information is derived from the categories’ sentence embeddings $\{t_{j}\}$, also contributing to $s_{ij}$. Quantitative information from object counts $\{num_{j}\}$ serves as constraints in (3). The problem is efficiently solvable using standard BIP solvers like PuLP mitchell2011pulp . The resulting non-zero entries in $X$ define the final set of recognized objects $\{(mask_{i},cat_{j})|x_{ij}=1\}$.
@@ 5.2 Results on EarthInstruct
Table 2 summarizes the results under open-ended setting. InstructSAM consistently achieves higher $\text{F}_{\text{1}}$-scores than remote sensing-specific methods, including those trained on grounded description tasks luo2024skysensegpt ; shabbir2025geopixel . Notably, InstructSAM surpasses LAE-Label pan2025locate by leveraging a global view of the image to accurately predict object categories. While the absence of class-specific instructions in this setting limits further gains, InstructSAM still demonstrates robust performance (Figure 4).
@@
Figure 4: Qualitative comparison of object detection and segmentation methods on DIOR dataset. InstructSAM consistently outperforms other methods in generating accurate bounding boxes and segmentation masks across various settings. More qualitative results can be found in Appendix G.
@@
Figure 5 compares inference time across open-ended methods equipped with 7B LLMs. InstructSAM exhibits nearly constant inference speed, in contrast to other approaches whose runtime increases linearly. Unlike representing bounding boxes as natural language tokens, InstructSAM reduces output tokens by 89% and total inference time by 32% compared to Qwen2.5-VL. This advantage will grow more pronounced as the model size scales up, highlighting the efficiency of our framework.
Table 4: Open vocabulary counting performance (Precision/Recall/$\text{F}_{\text{1}}$-score).
@@ 5.3 Ablation Studies
To assess InstructSAM’s generalization and scalability, we ablate LVLM counters, mask proposers, and CLIP models for the open-vocabulary detection (OVD) task on NWPU-VHR-10 (Table 5). InstructSAM consistently benefits from CLIP models fine-tuned on remote sensing data liu2024remoteclip ; wang2024skyscript over generic CLIP fang2024data , whereby yielding higher Box $\text{F}_{\text{1}}$-scores. Performance improves with larger model components, demonstrating the framework’s scalability. Notably, even with a smaller SAM2-S and SkyCLIP-B, InstructSAM coupled with Qwen2.5VL (40.6 Box $\text{F}_{\text{1}}$) outperforms direct detection using Qwen2.5VL alone (36.4 Box $\text{F}_{\text{1}}$), underscoring the efficacy of our approach.
@@
Using fixed thresholds to filter CLIP predictions cheng2024yolo bears inherent limitations. Figure 6 shows its performance is highly threshold-sensitive, consistent with the findings in li2025simple . Each category requires different optimal values, making a single threshold ineffective across classes. InstructSAM’s counting-constrained matching eliminates this dependency by dynamically adjusting assignments based on predicted counts, yielding superior results in multi-class and open-world scenarios.
@@
5.4 Error Analysis of OVD Task
@@ Appendix A Comparison with Related work
Method Publication Object Recognition Training Data Dense Segmentation OVD OED OSD
@@
SkysenseGPT luo2024skysensegpt aiXiv’24 FIT-RS
@@
Qwen2.5-VL bai2025qwen2 arXiv’25 Various Source
@@ C.3 Prompts in Open-Subclass Setting
[…]
@@
[…]
@@
[…]
@@
[…]
@@ D.3 Selection of Prompt Format
Appendix E Limitations
InstructSAM relies heavily on pre-trained foundation models. Consequently, its performance is limited by the capabilities and biases of these models. For example, SAM2 may struggle with segmenting the entire structure of objects with intricate geometries. Additionally, CLIP finetuned on scene-level remote sensing image-text pairs, performs suboptimally on aligning object-level image crops with semantic cues.
Future work could address these challenges by incorporating foundation models trained on more semantically diverse remote sensing data or leveraging advanced class-agnostic region proposers xie2025llama . Additionally, using InstructSAM as an automated labeling pipeline offers a promising avenue for constructing large-scale, semantically rich object recognition datasets.
Appendix F Broader Impacts
@@
Appendix G Qualitative Results
@@
G.1 Qualitative Results in Open-Vocabulary Setting
@@
G.2 Qualitative Results in Open-Ended Setting
@@
G.3 Qualitative Results in Open-Subclass Setting
@@
G.4 Qualitative Results on Other Instructions
@@
G.5 Qualitative Results in Natural Images

### v2

@@ Abstract
Language-guided object recognition in remote sensing imagery is crucial for large-scale mapping and automated data annotation. However, existing open-vocabulary and visual grounding methods rely on explicit category cues, limiting their ability to handle complex or implicit queries that require advanced reasoning. To address this issue, we introduce a new suite of tasks, including Instruction-Oriented Object Counting, Detection, and Segmentation (InstructCDS), covering open-vocabulary, open-ended, and open-subclass scenarios. We further present EarthInstruct, the first InstructCDS benchmark for earth observation. It is constructed from two diverse remote sensing datasets with varying spatial resolutions and annotation rules across 20 categories, necessitating models to interpret dataset-specific instructions. Given the scarcity of semantically rich labeled data in remote sensing, we propose InstructSAM, a training-free framework for instruction-driven object recognition. InstructSAM leverages large vision-language models to interpret user instructions and estimate object counts, employs SAM2 for mask proposal, and formulates mask-label assignment as a binary integer programming problem. By integrating semantic similarity with counting constraints, InstructSAM efficiently assigns categories to predicted masks without relying on confidence thresholds. Experiments demonstrate that InstructSAM matches or surpasses specialized baselines across multiple tasks while maintaining near-constant inference time regardless of object count, reducing output tokens by 89% and overall runtime by over 32% compared to direct generation approaches. We believe the contributions of the proposed tasks, benchmark, and effective approach will advance future research in developing versatile object recognition systems. The code is available at https://VoyagerXvoyagerx.github.io/InstructSAM/.
@@ 1 Introduction
Recent advancements in VLMs have demonstrated impressive performance in object detection liu_grounding_2023 ; lin2024generative ; pi2024ins ; pan2025enhance, semantic segmentation yu2023convolutions ; cho2024cat ; zhang2025mirsam, visual grounding bai2025qwen2 ; chen2024expanding and reasoning-based segmentation  ren2024pixellm ; lai2024lisa within the natural image domain. However, transferring these methods to remote sensing imagery presents several challenges. First, direct inference leads to significant accuracy degradation due to the domain gap between natural and aerial images fu2024cross . Second, most existing remote sensing open-vocabulary detection li2024toward ; li2024exploiting ; zang2024zero ; li2025semantic and segmentation zermatten2025learning ; ye2025towards ; huang2025zori methods are trained on datasets with only a limited number of categories, restricting their generalization to diverse unseen categories. Third, conventional detectors li2022grounded ; cheng2024yolo ; minderer2023scaling rely on a threshold to filter predicted bounding boxes, which is not obtainable in zero-shot scenarios.
@@ 2.2 Instruction-oriented remote sensing object detection and segmentation
Recent advances in VLMs radford2021learning ; liu2023visual have also enabled open-vocabulary learning in the remote sensing domain. Diverse semantic tags from OpenStreetMap haklay2008openstreetmap and those generated by LVLMs drive the development of contrastive language-image pre-training in remote sensing images zhang2024rs5m ; wang2024skyscript . Following the generic open-vocabulary learning frameworks, methods for remote sensing open-vocabulary detection li2024toward ; wei2024ova ; pan2025locate ; huang2025openrsd and segmentation zermatten2025learning ; ye2025towards ; huang2025zori ; li2024segearth methods emerge. However, their human-annotated training data, being limited to a few dozen categories li2020object ; waqas2019isaid ; sun2022fair1m ; liu2025control, hinder generalization to out-of-distribution or zero-shot scenarios. Although some remote sensing LVLMs could support object recognition tasks like single-class object detection kuckreja2024geochat ; luo2024skysensegpt ; irvin2024teochat , visual grounding muhtar2024lhrs ; pang2025vhm , referring expression segmentation zhou2024geoground ; ou2025geopix , grounded conversation generation shabbir2025geopixel , and scene graph generation luo2024skysensegpt , they failed to follow complex reasoning instructions, such as open-vocabulary and open-subclass object detection. To annotate vast-vocabulary training data for remote sensing object detection, LAE-Label pan2025locate employs a generic LVLM chen2024far to predict categories for cropped mask proposals. However, this approach loses global context for accurate category classification.
@@ 4.3 Mask-label matching with counting constraints
As depicted in Figure 3, this BIP formulation elegantly fuses visual information, semantic information, and quantitative information. The visual information comes from the CLIP embeddings of mask proposals $\{v_{i}\}$, which contribute to $s_{ij}$. Semantic information is derived from the categories’ sentence embeddings $\{t_{j}\}$, also contributing to $s_{ij}$. Quantitative information from object counts $\{num_{j}\}$ serves as constraints in (3). The problem is efficiently solvable using open-source BIP solvers like PuLP mitchell2011pulp . The resulting non-zero entries in $X$ define the final set of recognized objects $\{(mask_{i},cat_{j})|x_{ij}=1\}$.
@@ 5.2 Results on EarthInstruct
Table 2 summarizes the results under open-ended setting. InstructSAM consistently achieves equivalent or higher $\text{F}_{\text{1}}$-scores than remote sensing-specific methods, including those trained on grounded description tasks luo2024skysensegpt ; shabbir2025geopixel . Notably, InstructSAM surpasses LAE-Label pan2025locate by leveraging a global view of the image to accurately predict object categories. While the absence of class-specific instructions in this setting limits further gains, InstructSAM still demonstrates robust performance (Figure 4).
@@
Figure 4: Qualitative comparison of object detection and segmentation methods on DIOR dataset. InstructSAM consistently outperforms other methods in generating accurate bounding boxes and segmentation masks across various settings. More qualitative results can be found in Appendix F.
@@
Figure 5 compares inference times for open-ended methods using 7B LLMs. Mask proposal dominates the runtime of InstructSAM, whereas the PuLP optimizer solves the BIP problem in 0.07 seconds. InstructSAM maintains nearly constant inference time, while other methods scale linearly with the number of objects. Unlike representing bounding boxes as natural language tokens, InstructSAM reduces output tokens by 89% and total inference time by 32% compared to Qwen2.5-VL. This advantage will grow more pronounced as the model size scales up, highlighting the efficiency of our framework.
[…]
@@ 5.3 Ablation studies
To assess InstructSAM’s generalization and scalability, we ablate LVLM counters, mask proposers, and CLIP models for the open-vocabulary detection (OVD) task on NWPU-VHR-10 (Table 5). InstructSAM consistently benefits from CLIP models fine-tuned on remote sensing data liu2024remoteclip ; wang2024skyscript over generic CLIP fang2024data , whereby yielding higher Box $\text{F}_{\text{1}}$-scores. Performance improves with larger model components, demonstrating the framework’s scalability. Notably, even with a smaller SAM2-S and SkyCLIP-B, InstructSAM coupled with Qwen2.5-VL (40.6 Box $\text{F}_{\text{1}}$) outperforms direct detection using Qwen2.5-VL alone (36.4 Box $\text{F}_{\text{1}}$), underscoring the efficacy of our approach.
@@
Using fixed thresholds to filter CLIP predictions cheng2024yolo has inherent limitations. In Figure 6, the green and blue solid curves show InstructSAM without matching, where predictions are filtered by a similarity score (CLIP score) threshold before evaluation. Optimal thresholds vary widely across categories. As a result, a single global value performs poorly, and the results are highly threshold-sensitive, which is consistent with the findings in  li2025simple . In contrast, InstructSAM’s counting‑constrained matching removes this dependency by dynamically assigning predictions according to the estimated counts, yielding stronger performance in multi‑class and open‑world settings.
5.4 Error analysis of OVD task
@@
6 Limitations
InstructSAM builds on pre-trained foundation models, and its performance inherits their capabilities and biases. For instance, SAM2 can miss the full extent of objects with intricate geometry (e.g., basketball court in Figure 4). CLIP models finetuned on scene-level remote sensing image–text pairs perform suboptimally when aligning object-level crops with semantic cues. More generally, foundation models such as GPT‑4o and SAM are trained primarily on optical imagery and struggle to count or segment objects in synthetic aperture radar (SAR) images.
Future work can mitigate these limitations by (i) incorporating foundation models trained on semantically diverse, multi-modal remote sensing data and (ii) leveraging advanced class-agnostic region proposers xie2025llama . Extending InstructSAM to semantic segmentation would further benefit the earth observation community. Although InstructSAM is only a first step, we aim to use it as an automated labeling engine to annotate at scale, enabling large, semantically rich object-recognition datasets.
@@ 7 Conclusions
Acknowledgement
This work was supported by the National Earth Observation Data Center Research Project, National Natural Science Foundation of China (62506229), and Natural Science Foundation of Shanghai (25ZR1402268).
@@ Appendix A Comparison with related work
Method Publication
Object Recognition
Training Data
Dense
Segmentation
OVD OED OSD
@@
SkysenseGPT luo2024skysensegpt arXiv’24 FIT-RS
@@
Qwen2.5-VL bai2025qwen2 arXiv’25 Various sources
@@ C.3 Prompts in open-subclass setting
    "If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
@@
    "If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
@@
    "Return only the categories and counts that meet the visibility and resolution criteria. If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}".
@@
    "Return only the categories and counts that meet the visibility and resolution criteria. If none of the objects belong to the parent category is visible, output a empty JSON object like $\{$ \}"
@@ D.3 Selection of prompt format
Appendix E Broader impacts
@@
Appendix F Qualitative results
@@
F.1 Qualitative results in open-vocabulary setting
@@
F.2 Qualitative results in open-ended setting
@@
F.3 Qualitative results in open-subclass setting
@@
F.4 Qualitative results on other instructions
@@
F.5 Qualitative results in natural images